OpenStudy (anonymous):

Is the square root of negative 4/9 a positive or negative number?

OpenStudy (anonymous):

neither. imaginary numbers do not get an adjective "positive" or "negative"

OpenStudy (anonymous):

so what do i put as my answer? 2/3 is my answer but when square rooted from 4/9 there is a negative sign in front of it. so whats my answer?

jimthompson5910 (jim_thompson5910):

\[-\sqrt{\frac{4}{9}}=-\frac{2}{3}\]

jimthompson5910 (jim_thompson5910):

based off the last problem, this is what I'm guessing the expression is

OpenStudy (anonymous):

i read it is \[\sqrt{-\frac{9}{4}}\]

OpenStudy (anonymous):

which will give you \[\frac{3}{2}i\] neither positive nor negative

OpenStudy (anonymous):

yes jim. thats the problem. 

OpenStudy (anonymous):

really depends on what it looks like originally. you wrote the square root of negative 4/9

OpenStudy (anonymous):

ooooooooooh ok ignore me. jim has it

jimthompson5910 (jim_thompson5910):

ok thx, in the future to say \[-\sqrt{\frac{4}{9}}\] say "negative square root of 4/9"

OpenStudy (anonymous):

sorry. i dont have the symbols on my keyboard

jimthompson5910 (jim_thompson5910):

since "square root of negative 4/9" means \[\large \sqrt{-\frac{4}{9}}\]

OpenStudy (anonymous):

or "minus the square root of 4/9" put the "minus" or "negative" before the root

OpenStudy (anonymous):

it would be a complex number, you can't square root negatives so you have to bring in imaginary numbers: \[\sqrt{-4/9}\] = \[\sqrt{-4}\div \sqrt{9}\] make the \[\sqrt{-4} = \sqrt{-1}*\sqrt{4} = 2i\] so \[2i/\sqrt{9}\] thats 2i/3 - a complex number

jimthompson5910 (jim_thompson5910):

probably best to just draw it out next time since that worked

OpenStudy (anonymous):

thank you jim. your answer worked best for me.

jimthompson5910 (jim_thompson5910):

yw